OpenStudy (anonymous):

The length of a rectangular field is 4 m longer than its width. If the perimeter is 56 m, what is the width of the field?

OpenStudy (anonymous):

l=w+4

OpenStudy (anonymous):

srry im a little tired...

OpenStudy (anonymous):

to find the perimeter, u need to add 2 length and 2 width

OpenStudy (anonymous):

so 2w+2(w+4)=56

OpenStudy (anonymous):

2w+2w+8=56 2w+2w=56-8=48 4w=48 w=48/4 w=12

OpenStudy (anonymous):

l=w+4 l=12+4 l=16

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np